# Synthetic SQL Column Descriptions and Their Impact on Text-to-SQL Performance

### v3

@@ Abstract
Relational databases often suffer from uninformative descriptors of table contents, such as ambiguous columns and hard-to-interpret values, impacting both human users and text-to-SQL models. In this paper, we explore the use of large language models (LLMs) to automatically generate detailed natural language descriptions for SQL database columns, aiming to improve text-to-SQL performance and automate metadata creation. We create a dataset of gold column descriptions based on the BIRD-Bench benchmark, manually refining its column descriptions and creating a taxonomy for categorizing column difficulty. Through evaluating several LLMs, we find that incorporating these column descriptions consistently enhances text-to-SQL model performance, particularly for larger models like GPT-4o, Qwen2 72B and Mixtral 22Bx8. However, models struggle with columns that exhibit inherent ambiguity, highlighting the need for manual expert input. Notably, Qwen2-generated descriptions, containing by annotators deemed superfluous information, outperform manually curated gold descriptions, suggesting that models benefit from more detailed metadata than humans expect. Future work will investigate the specific features of these high-performing descriptions and explore other types of metadata, such as numerical reasoning and synonyms, to further improve text-to-SQL systems. The dataset, annotations and code will all be made available.
@@ 1 Introduction
The recent advancements in LLMs have shown that they can produce coherent and contextually relevant text across various domains [7], making them potential tools for automating such metadata creation. In the context of SQL databases, the hypothesis is that LLMs could leverage multiple signals from a database to generate meaningful column descriptions, including the schema structure, relationships between tables, and the values present within the database itself. By analyzing these signals, LLMs can go beyond simple type detection, attempting to infer a column’s broader purpose and a more detailed explanation. Previous research has used generative LLMs to expand abbreviated column names in a SQL setting [8], but to the best of our knowledge, no works exist for generating more detailed descriptions across a wider array of column types and formats.
@@
  • We create a new dataset based on the BIRD-Bench benchmark, improving the quality of column descriptions through a combination of LLM-generated outputs and manual refinements. This dataset provides a strong foundation for evaluating column description generation.
@@ 2.2 Metadata Usage in Text-to-SQL
In contrast to these approaches, our work focuses on generating comprehensive SQL column descriptions directly using LLMs. While prior research emphasizes leveraging external knowledge and complex data retrieval processes, we aim to enhance the interpretability and usability of SQL queries by creating detailed descriptions for SQL columns.
@@ 4.3 Text-to-SQL Evaluations
The experimental setup involved five scenarios: (1) Without any column descriptions, where the model operates with only the database schema; (2) With BIRD-Bench original column descriptions, utilizing the original descriptions provided by BIRD-Bench in addition to the schema; (3) With gold standard descriptions, where the model uses our manually refined gold standard descriptions in addition to the schema; (4) With GPT-4o generated column descriptions, where the model uses descriptions generated by the best model from Section 4.1 in addition to the schema; and (5) With Qwen2 generated descriptions, where the model uses descriptions rated with a high amount of Superflous information. We included both the GPT-4o and the Qwen2-generated descriptions to test whether our humanly defined categories of "Perfect" and "Superfluous information" are aligned with what is practically useful for LLMs in the text-to-SQL task.
In all scenarios, we included only column descriptions as the available metadata alongside the database schema. We set the model temperature to 0. The full prompt for the text-to-SQL generation is presented in Appendix B.9.
5 Results
@@ 6 Conclusions and Future Work
In this work, we demonstrated that natural language descriptions and metadata are crucial for improving the performance of text-to-SQL models. Our experiments show that including detailed column descriptions, whether manually refined or synthetically generated, consistently enhances query generation accuracy across a range of LLMs. However, generating accurate descriptions for columns with inherent ambiguity remains a challenge, as models unsurprisingly struggle to handle such columns effectively.

### v4

@@ Abstract
Relational databases often suffer from uninformative descriptors of table contents, such as ambiguous columns and hard-to-interpret values, impacting both human users and text-to-SQL models. In this paper, we explore the use of large language models (LLMs) to automatically generate detailed natural language descriptions for SQL database columns, aiming to improve text-to-SQL performance and automate metadata creation. We create a dataset of gold column descriptions based on the BIRD-Bench benchmark, manually refining its column descriptions and creating a taxonomy for categorizing column difficulty. We then evaluate several different LLMs in generating column descriptions across the columns and different difficulties in the dataset, finding that models unsurprisingly struggle with columns that exhibit inherent ambiguity, highlighting the need for manual expert input. We also find that incorporating such generated column descriptions consistently enhances text-to-SQL model performance, particularly for larger models like GPT-4o, Qwen2 72B and Mixtral 22Bx8. Notably, Qwen2-generated descriptions, containing by annotators deemed superfluous information, outperform manually curated gold descriptions, suggesting that models benefit from more detailed metadata than humans expect. Future work will investigate the specific features of these high-performing descriptions and explore other types of metadata, such as numerical reasoning and synonyms, to further improve text-to-SQL systems. The dataset, annotations and code will all be made available.
@@ 1 Introduction
The recent advancements in LLMs have shown that they can produce coherent and contextually relevant text across various domains [7], making them potential tools for automating such metadata creation. In the context of SQL databases, the hypothesis is that LLMs could leverage multiple signals from a database to generate meaningful column descriptions, including the schema structure, relationships between tables, and the values present within the database itself. Previous research has used generative LLMs to expand abbreviated column names in a SQL setting [8], but to the best of our knowledge, no works exist for generating more detailed descriptions across a wider array of column types and formats.
@@
  • We create a new dataset of column descriptions based on the BIRD-Bench benchmark, improving the quality of existing column descriptions through a combination of LLM-generated outputs and manual refinements.
@@ 2.2 Metadata Usage in Text-to-SQL
In contrast to these approaches, our work focuses on generating comprehensive SQL column descriptions directly using LLMs, rather than using humanly curated descriptions.
@@ 4.3 Text-to-SQL Evaluations
The experimental setup involved five scenarios: (1) Without any column descriptions, where the model operates with only the database schema; (2) With BIRD-Bench original column descriptions, utilizing the original descriptions provided by BIRD-Bench in addition to the schema; (3) With gold standard descriptions, where the model uses our manually refined gold standard descriptions in addition to the schema; (4) With GPT-4o generated column descriptions, where the model uses descriptions generated by the best model from Section 4.1 in addition to the schema; and (5) With Qwen2 generated descriptions, where the model uses descriptions rated with a high amount of Superflous information. We included both the GPT-4o and the Qwen2-generated descriptions to test whether our humanly defined categories of "Perfect" and "Superfluous information" are aligned with what is practically useful for LLMs in the text-to-SQL task. In all scenarios, we included only column descriptions as the available metadata alongside the database schema. We set the model temperature to 0. The full prompt for the text-to-SQL generation is presented in Appendix B.9.
5 Results
@@ 6 Conclusions and Future Work
In this work, we demonstrated that natural language descriptions and metadata are crucial for improving the performance of text-to-SQL models. Our experiments show that including detailed column descriptions, whether manually refined or synthetically generated, consistently enhances query generation accuracy across a range of LLMs. However, generating accurate descriptions for columns with inherent ambiguity unsurprisingly remains a challenge. For handling such columns we propose a two-step approach: first, automatically generate descriptions for columns with clear interpretation signals in the database; then, have a domain expert manually describe inherently ambiguous columns. This ensures LLMs handle straightforward columns independently while relying on expert input for ambiguous cases, though further research is needed to pinpoint these columns.
@@
Acknowledgements
The computations in this paper was enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre in Linköping, Sweden, to whom we extend our gratitude. Additionally, this research was also partially funded by ELLIIT, a strategic research environment in information technology and mobile communications funded by the Swedish government.